# Reasoning Efficiently Through Adaptive Chain-of-Thought Compression: A Self-Optimizing Framework

### v1

@@ 7.2.2. External Validity
External validity concerns the generalizability of our findings beyond the datasets and settings considered in this work. Our experiments primarily focus on software engineering tasks, as they represent one of the most critical application domains for reasoning models. To mitigate the risk of over-specialization, we additionally included a mathematical reasoning dataset (GSM-8K), which provides a different reasoning style and serves as an initial step toward broader validation. Nevertheless, important domains such as commonsense reasoning, legal QA, and scientific problem solving remain unexplored.
Moreover, all evaluations were conducted with fixed prompts and prompting strategies. While our method is designed to be prompt-agnostic, alternative instruction formats, few-shot configurations, or domain-specific prompts could influence the observed trade-off between compression and accuracy. Finally, although the datasets used contain tasks of varying difficulty, we have not systematically examined whether compression benefits differ across problem complexity levels.
To further strengthen external validity, we plan to expand our evaluation to additional reasoning domains and more diverse datasets, as well as investigate robustness under alternative prompting strategies. Future work will also include stratified analyses by task difficulty to better understand whether SEER is equally effective on both simple and challenging instances.
@@ 8. Related Works
$\blacksquare$ Prompt-Guided Efficient Reasoning. Prompt-based approaches improve reasoning efficiency by explicitly guiding the model to be concise. CCoT (Cheng and Van Durme, 2024) uses a concise CoT template to encourage brevity through prompt wording. Token-Budget (Han et al., 2024) proposes a token-budget-aware reasoning framework that dynamically adjusts reasoning length based on estimated complexity. CoD (Xu et al., 2025a) limits reasoning to within five words via a “Chain-of-Draft” template. Token Complexity (Lee et al., 2025) systematically analyzes prompt compression strategies, highlighting that many still operate far from theoretical efficiency limits.
@@
$\blacksquare$ Chain-of-Thought Effectiveness in Software Engineering. CoT has shown strong benefits in software engineering tasks. SCoT (Li et al., 2025) was among the first to apply structured CoT to code generation. The emergence of Deepseek-R1 (DeepSeek-AI, 2025b) further highlighted the value of streamlined CoT for software engineering tasks. Liu et al. (Liu et al., 2025) revisited CoT training paradigms, showing that some seemingly effective methods do not transfer well to real-world coding tasks. UnCert-CoT (Zhu et al., 2025) uses uncertainty estimation to decide when to trigger CoT, avoiding unnecessary reasoning. CoD (Yang, 2025) adapts the “Chain-of-Draft” approach to software engineering tasks, greatly reducing token usage.
However, existing SE-focused methods often target specific task scenarios or rely on inference-time control, which may not eliminate pathological failure modes, such as infinite reasoning loops. Our method differs by explicitly filtering out redundant or looping reasoning during training, allowing the model to internalize concise, stable reasoning behaviors. This makes it robust not only in software engineering tasks but also in mathematical reasoning, while retaining interpretability and avoiding the need for task-specific prompt engineering.
@@ 9. Conclusion
Our work systematically examined the inefficiency of excessively long CoT reasoning, identifying that long CoT can lead to truncation and overthinking, resulting in decreased accuracy. We then present SEER, a self-enhancing framework for adaptive CoT compression. Through extensive experiments on software engineering and mathematical reasoning tasks, SEER reduces CoT length by an average of 42.1% while maintaining or even improving accuracy. Our ablation results confirm the complementary roles of BoN sampling and adaptive filtering, and cross-domain evaluations demonstrate SEER’s strong generalizability. SEER also effectively mitigates truncation and suppresses infinite reasoning loops, addressing key efficiency and robustness issues. These findings highlight SEER as a practical approach for making CoT-enhanced LLMs more efficient and reliable. Future work will extend SEER to broader reasoning domains, apply it to diverse model architectures, and explore integration into latency-sensitive real-world systems.

### v2

@@ 7.2.2. External Validity
External validity concerns whether our findings generalize beyond the evaluated settings. A potential limitation is that our study focuses on software engineering tasks, which may restrict applicability to other reasoning domains. We mitigate this threat by evaluating SEER on three representative software engineering benchmarks covering distinct task types, including code generation, defect detection, and code search, each involving different reasoning patterns.
Another concern is sensitivity to prompt design. To reduce this risk, all methods are evaluated using fixed prompts and consistent prompting strategies, ensuring that observed improvements are attributable to the proposed method rather than prompt-specific effects. Moreover, SEER is designed to be largely prompt-agnostic, further limiting reliance on particular instruction formats.
Finally, the effectiveness of CoT compression may vary with problem difficulty. Although we do not explicitly stratify instances by complexity, the evaluated datasets naturally contain problems of varying difficulty, and results are reported over full benchmarks rather than curated subsets. A finer-grained analysis by difficulty is left for future work.
@@ 8. Related Works
$\blacksquare$ Prompt-Guided Efficient Reasoning. Prompt-based approaches improve reasoning efficiency by explicitly guiding the model to be concise. CCoT (Cheng and Van Durme, 2024) uses a concise CoT template to encourage brevity through prompt wording. Token-Budget (Han et al., 2024) proposes a token-budget-aware reasoning framework that dynamically adjusts reasoning length based on estimated complexity. CoD (Xu et al., 2025b) limits reasoning to within five words via a “Chain-of-Draft” template. Token Complexity (Lee et al., 2025) systematically analyzes prompt compression strategies, highlighting that many still operate far from theoretical efficiency limits.
@@
$\blacksquare$ Chain-of-Thought Effectiveness in Software Engineering. CoT has shown strong benefits in software engineering tasks. SCoT (Li et al., 2025) was among the first to apply structured CoT to code generation. The emergence of DeepSeek-R1 (DeepSeek-AI, 2025) further highlighted the value of streamlined CoT for software engineering tasks. Liu et al. (Liu et al., 2025) revisited CoT training paradigms, showing that some seemingly effective methods do not transfer well to real-world coding tasks. UnCert-CoT (Zhu et al., 2025) uses uncertainty estimation to decide when to trigger CoT, avoiding unnecessary reasoning. CoD (Yang, 2025) adapts the “Chain-of-Draft” approach to software engineering tasks, greatly reducing token usage.
However, existing SE-focused methods often target specific task scenarios or rely on inference-time control, which may fail to eliminate severe failure modes such as looping behaviors. Our method differs by explicitly filtering out redundant or looping reasoning during training, allowing the model to internalize concise, stable reasoning behaviors. This makes it robust in software engineering tasks while retaining interpretability and avoiding the need for task-specific prompt engineering.
@@ 9. Conclusion
In this paper, we investigate the practical costs of overly long CoT reasoning in software engineering settings. We find that many modern reasoning models produce excessively verbose CoTs, which substantially increases token cost and can trigger truncation; moreover, many truncations are caused by looping behaviors. To address these challenges, we proposed SEER, a self-enhancing framework for adaptive CoT compression that learns concise reasoning patterns from self-generated outputs by combining BoN sampling with a lightweight, data-driven CoT length filter. Across three software engineering tasks, SEER reduces CoT length by an average of 41.6% while improving performance. Cross-domain evaluations indicate good generalization to unseen benchmarks. Overall, SEER makes CoT-enhanced LLMs more efficient and robust under real-world context constraints. Future work includes extending SEER to more domains and model families and integrating it into latency-sensitive agent pipelines.